OpenStudy (anonymous):

Does anyone know how to write this in sigma notation? -2, + 1 + 4 + ... + 13 Please explain(:

OpenStudy (anonymous):

So its -2+(-2+3*1)+(-2+3*2)+(-2+3*3)+...an

OpenStudy (anonymous):

so adding 3... Where do we go from there?(:

OpenStudy (anonymous):

Can you come up with an expression for the n'th term of the sequence that is being added here. -2,(-2+3*1),(-2+3*2),(-2+3*3),...an

OpenStudy (anonymous):

plug in for this formula tn= t1 +(n-1)d ?????

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

tn = -2 + (n-1) 3 so -2 + 3n - 3 -5 + 3n What do we do next to put it in sigma?

OpenStudy (anonymous):

Over what values of n are you adding. What is the n for the first value and for the last value?

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

No what value of n gives you 13 from your tn equation. Also which one gives you -2?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

You have tn=-5 + 3n. Now set tn=-2 and solve for n.

geerky42 (geerky42):

\[\Large \sum_{n=0}^{5} 3n-2\]

OpenStudy (anonymous):

wait why do we set tn = -2

OpenStudy (anonymous):

help(:

OpenStudy (anonymous):

geerky that isn't the right answer...